# Literature Filtering for Systematic Reviews with Transformers

### v1

@@
Literature Filtering for Systematic Reviews with Transformers
John Hawkins $0$-0001-6507-3671 Transitional AI Research GroupSydenyAustralia43017-6221  and  David Tivey Royal Australasian College of Surgeons (RACS)Australia
Systematic Review, Literature Screening, Machine Learning, BERT, Transfer Learning, Natural Language Processing
1. Abstract
@@
Identifying critical research within the growing body of academic work is an essential element of quality research. Systematic review processes, used in evidence-based medicine, formalise this as a procedure that must be followed in a research program. However, it comes with an increasing burden in terms of the time required to identify the important articles of research for a given topic. In this work, we develop a method for building a general-purpose filtering system that matches a research question, posed as a natural language description of the required content, against a candidate set of articles obtained via the application of broad search terms. Our results demonstrate that transformer models, pre-trained on biomedical literature then fine tuned for the specific task, offer a promising solution to this problem. The model can remove large volumes of irrelevant articles for most research questions.
2. Introduction
@@
Systematic Reviews are an essential aspect of evidence-based medicine. They ensure that researchers are aware of the most relevant literature on a particular topic. They are also time-consuming, error-prone and costly to undertake(Dinter2021b, ). The expanding volumes of literature and growing number of active research areas are increasing the burden of thorough systematic reviews(Bastian2010, ; Miwa2014, ).
The problem of matching a research question to a set of articles sits within the broader problem of matching queries to documents, which is plagued by the issue of algorithmically identifying what is semantically relevant between a query and document pair(Ostendorff2020, ). A common approach to building software applications that assist the systematic review process focuses on tools that refine database literature searches and results, including the use of empirical methods for refining search terms (Wilczynski2004, ; Zdravevski2019, ; Weier2020, ), automated query generation(Scells2020, ), and recommendation methods based on the researcher’s publication profile(Chen2016, ) and machine learning models that emulated the identification of articles of clinical importance(Aphinyanaphongs2005, ).
@@
Recent development has focused on methods to improve the efficiency of the initial literature screening, including processes that reduce the human workload through question-specific predictive models built on an initial set of human reviewer labels(Frunza2010, ; Bannach-Brown2019, ), expanded to include other kinds of meta-data(Olorisade2019, ) and applications that update the prioritisation of abstracts in an iterative feedback process (Jonnalagadda2013, ; Miwa2014, ; Kontonatsios2020, ; Bravo2021, ), or expansion of the articles for consideration using a similarity space and initial human selected abstracts (Pham2021, ). Recent machine learning studies for automation of systematic reviews focus on alternative machine learning approaches to building these task/question specific models, with Convolutional Neural Networks (CNNs) (Dinter2021a, ), and Transformer Networks (Aum2021, ). Multiple studies are making incremental steps toward more general models that reuse what has been learned from previous systematic reviews in different kinds of transfer learning. This includes learning document embeddings that capture information relevant to literature filtering(Kontonatsios2020, ), and clustering candidate article abstracts with the articles from previous literature reviews to identify high-quality candidates(Cawley2020, ).
@@
However, there is some question as to whether these tools provide any real value to systematic review teams(Gartlehner2019, ). The tools Abstrackr(Wallace2012, ) and EPPI-Reviewer(Thomas2010, ) can both be trained on custom datasets to filter articles toward a subset for human reviewers. Independent evaluation has indicated that these tools will reduce the burden on human reviewers by a maximum of 50-60%, respectively. However, on some research questions, they perform considerably worse(Tsou2020, ). Arguably, the biggest challenge for this approach is the requirement for an initial set of labelled data for the specific question that is large enough to provide a viable model(Aum2021, ).
@@
Alternative approaches use specialised, but general-purpose models, to identify factors like the presence of randomised control trials within a given study(Cohen2015, ), or the specific text sections that correspond to the elements of the PICO(Richardson1995, ) framework (population; intervention; comparator; outcomes) (Wallace2016, ; Kang2019, ; Jin2020, ). Multiple factors about a paper can be used to filter articles in a systematic review, including the scientific rigor of the content (Ambalavanan2020, ), the quality of the research irrespective of the content(Lokker2023, ). There are multiple datasets available for these tasks(Nye2018, ; DeYoung2020, ; Ambalavanan2020, ), and multiple modelling approaches applied to this task with a recent focus on the use of neural networks from Long Short Term Memory (LSTM) (Jin2018, ) to Transformer models(Schmidt2020, ; Ambalavanan2020, ).
@@
Review articles from the systematic review literature itself have concluded that machine learning based filters that reduce the volume of articles needed to be reviewed are the most promising line of research in the automation of systematic reviews (Marshall2019, ; OMara-Eve2015, ).
@@
The Transformer model(Vaswani2017, ) has proven to be an effective tool for building text-based classification systems across a range of domains. Part of the utility comes from the fact that it can be pre-trained on the language from a domain and then used for specific tasks, allowing the previous learning to be ’transferred’ and applied. It has been shown that transformer models pre-trained on academic abstracts can be used to create state-of-the-art classifiers for systematic review systems, but only when they train the models on data for a specific task (Aum2021, ), extending previous work with feed-forward neural networks to learn a document embedding that can be employed via transfer learning for a specific systematic review task(Kontonatsios2020, ). The requirement for task-specific training places a burden on teams in terms of the time and human effort required to provide labels for the initial classifier. The open research question is whether we can use modern natural language processing transformers to develop general-purpose filters that learn the general correspondence between the phrasing of a research question and the relevant articles.
@@
This research outlines a dataset and comparison of methods for a general-purpose systematic review model. A single model is trained that is given input data combining a research question and information about a specific article. It then produces a score indicating the likelihood that the specific article would be included by a human reviewer during a literature pre-screen. Our results demonstrate that using BioBERT(Lee2020, ) version of the BERT Transformer(Devlin2019, ) pre-trained on biomedical literature can contribute a promising solution to this problem, but the ideal approach depends on the intended process for using a model to improve the systematic review process. We discuss the observed relationships between research questions and machine learning approaches, and conclude that more work is needed to determine the ideal model for all systematic reviews.
3. Methodology
@@
Figure 1. Schematic representation of the standard machine learning approach to building a systematic review filter tool. 1. Research question is proposed, 2. Database search is conducted, 3. List of candidate articles is compiled, 4. Human annotators review a small selection of articles, 5. Small annotated collection is used to train a machine learning filter, 6. The machine learning is applied to the remaining documents, 7. Producing the final set for complete review by human experts.
The standard approach to building systematic review tools with machine learning shown in Figure 1. The standard model requires an initial labelled set to seed the filter model.
@@
Figure 2. Schematic representation of the general systematic review filter process. The first five steps define how the general filter is trained, the last two the application to new research questions. 1. A set of training research question are taken, 2. Database searches are conducted for each, 3. Lists of candidate articles are compiled, 4. Human annotators review these training questions to create an annotated set, 5. The general filter is trained. 6. The machine learning is applied to filter documents for a new query, 7. Producing the final set for complete review by human experts.
@@
Our alternative approach is depicted schematically in Figure 2. The model is trained on a new dataset derived from multiple research questions with human-annotated lists of articles returned from a database search. The bottom frame demonstrates that this general purpose filter can now be applied to future research questions without the need for human annotators to provide the initial labeled set. Our approach learns an abstract relationship between research questions and articles that transcends individual research questions.
3.1. Dataset Curation
Table 1. Medical Research Questions Datasets
@@ 3.1. Dataset Curation
We restrict our initial analysis to research questions that are concerned with Therapeutics, a constrained domain within the scope of medical research topics. We identified 10 research questions from the ASERNIP-S systematic review archive for which a team of human researchers had manually labelled thousands of articles as either relevant or not relevant. These initial questions are summarised in Table 1, showing the total number of abstracts retrieved from bibliographic database (e.g, PubMed) searches and the number deemed relevant by the human reviewers.
@@
In addition to the natural language expression of the overall research question, we also generated natural language expressions for each of the four PICO elements independently. The PICO framework has been shown to be well suited to representing therapy questions(Huang2006, ), and we conduct experiments to determine whether text based machine learning tools are able to exploit this alternative framing of the research question.
@@
3.2. Machine Learning Models
Table 2. Text Match Feature Functions
@@ 3.2. Machine Learning Models
We approach this problem using four general machine learning strategies. The first strategy is a baseline approach using a set of standard feature engineering approaches and general-purpose machine learning algorithms ranging from simple to complex. In this approach the research question is contained in a single text field of the data and compared with the article’s title and abstract using text matching algorithms. Each of these text matching approaches generates a numeric value which either describes the context of the research article (title and abstract) or indicates the similarity of the text in the question and the article (title and abstract). The text features and matching approaches are drawn from multiple python packages and compiled into the texturizer python package (texturizer, ); the full set of feature generating functions we use are summarised in Table 2. Note, that the TFIDF approach is only used in some of the experiments, and in all cases this is indicated in the descriptions of the results. The majority of the algorithms used are drawn from the SciKit Learn package (full details of which can be found in the code base).
@@
In the third approach, we employ a set of pre-trained BERT Transformer models (Devlin2019, ) to generate sentence similarity scores as new features. We continue to use the standard text match features from strategies one and two, but simply append the BERT similarity features to the dataset. This approach is done using both the standard natural language representation of the research question and the PICO framework text. We use two general versions of BERT, the BioBERT model(Lee2020, ) and the BlueBERT model(peng2019, ). Summary details for these pre-trained BERT models is shown in Table 3, including the number of heads, layers and the vocabulary size of each model. These parameters provide an approximation to the complexity of each model.
@@
Table 3. Pre-Trained Transformers
@@
3.3. Experimental Design
We aim to design a system that can identify relevant abstracts based on a new research question proposed for systematic review. The ideal system cannot rely on a bootstrap set of initial labelled examples for the specific research question. To align our testing protocol with this goal we use a variation of the cross-validation methodology in which each fold of the cross-validation consists of all data for one research question. This Leave-One-Question-Out (LOQO) method provides a close approximation to real-world usage, we evaluate how well the model can identify the relevant articles based purely on the correspondence between the content of the research question and the content of the article’s title and abstract.
3.4. Evaluation Metrics
To evaluate our results we use multiple different metrics. The Area Under the Curve ($AUC$)(Hanley1982, ) provides a value bounded between zero and one that captures the overall ability of a model to rank all articles(Melo2013, ). In addition, we use two metrics to evaluate the performance of the model at the upper and lower end of the article ranking, each of which cater to specific use cases of the models. We first consider the use case of eliminating irrelevant candidate articles (a dominant method by which machine learning filtering tools can provide efficiency for systematic review processes(Marshall2019, ; OMara-Eve2015, )). To evaluate models for this application we calculate the accuracy in the bottom 50% ($AccBot50\%$). We take the 50% lowest scoring records for a research question and then calculate the percentage of them that were annotated as irrelevant by the human reviewers. The calculation is shown in Equation 3.1:
@@ 3.4. Evaluation Metrics
4. Results
We show the results for both standard machine learning approaches and transformer models in Table 4 and Table 5. These table display the $AUC$, and $AccBot50\%$, respectively. Each modelling approach is displayed as a row with the results for each of the research questions in the corresponding column. We have highlighted in yellow the best (or equal best) score for each research question.
Table 4. AUC by Model and Research Question

### v2

# Efficient Systematic Reviews: Literature Filtering with Transformers & Transfer Learning

@@
Efficient Systematic Reviews: Literature Filtering with Transformers & Transfer Learning
John Hawkins
Transitional AI Research Group
David Tivey
Royal Australasian College of Surgeons (RACS)
1 Abstract
@@
Identifying critical research within the growing body of academic work is an intrinsic aspect of conducting quality research. Systematic review processes used in evidence-based medicine formalise this as a procedure that must be followed in a research program. However, it comes with an increasing burden in terms of the time required to identify the important articles of research for a given topic. In this work, we develop a method for building a general-purpose filtering system that matches a research question, posed as a natural language description of the required content, against a candidate set of articles obtained via the application of broad search terms. Our results demonstrate that transformer models, pre-trained on biomedical literature, and then fine tuned for the specific task, offer a promising solution to this problem. The model can remove large volumes of irrelevant articles for most research questions. Furthermore, analysis of the specific research questions in our training data suggest natural avenues for further improvement.
Keywords— Systematic Review, Literature Screening, Machine Learning, BERT, Transfer Learning, Natural Language Processing
2 Introduction
@@
Systematic Reviews are an essential aspect of evidence-based medicine. They ensure that researchers are aware of the most relevant and up-to-date literature on a particular question. They are, however, labour intensive, time-consuming, error-prone and costly to undertake[1]. The expanding volumes of literature and growing active research areas are increasing the burden of thorough systematic reviews [2, 3]. Furthermore, it has been argued that the semantic complexity of scientific literature can hinder the process of systematic reviews, impeding both efficiency and discovery[4].
These issues are affected by a broader set of problems related to the computational management of academic literature and the curation of datasets of research outcomes. This includes matching references between datasets or from bibliographies to canonical records [5], and identifying the works written by a particular researcher. The problem of matching a research question to a set of articles sits within the broader problem of matching queries to documents, which is plagued by the issue of algorithmically identifying what is semantically relevant between a query and document pair[6]. The systematic review process is one of the more difficult problems in this domain, as it requires algorithms that can anticipate what is relevant to a very specific research question.
@@
There have been many efforts to build general software applications that assist in the systematic review process. Initial approaches focused on algorithms and tools that would improve and refine database literature searches, including the use of empirical studies into methods for refining search terms and strategies [7, 8, 9], automated query generation[10], and machine learning models that emulated the identification of articles of clinical importance[11]. Most focus has been on methods to improve the efficiency of the initial literature screening, labelling articles as either relevant or irrelevant based on title and abstract alone. This includes processes that reduce the human workload through question-specific predictive models built on an initial set of human reviewer labels[12, 13], expanded to include other kinds of meta-data[14] and applications that update the prioritisation of abstracts in an iterative feedback process [15, 3, 16], or expansion of the articles for consideration using a similarity space and initial human selected abstracts [17]. Recent machine learning studies for automation of systematic review focus on alternative machine learning approaches to building task/question specific models, including the use of Convolutional Neural Networks (CNNs) [18], and Transformer Networks [19]. Multiple studies are making incremental steps toward more general models that reuse what has been learned from previous systematic reviews in different kinds of transfer learning. This includes learning document embeddings that capture information relevant to literature filtering[16], and clustering candidate article abstracts with the articles from previous literature reviews to identify high-quality candidates[20].
@@
However, there is some question as to whether these tools provide any real value to systematic review teams[21]. The tools Abstrackr[22] and EPPI-Reviewer[23] can both be trained on custom datasets to filter articles toward a subset for human reviewers. Independent evaluation has indicated that these tools will reduce the burden on human reviewers by a maximum of 50-60%, respectively. However, on some research questions, they perform considerably worse[24]. Arguably, the biggest challenge for this approach is the requirement for an initial set of labelled data for the specific question that is large enough to provide a viable model[19].
@@
There are alternative approaches that use specialised, but general-purpose models, to assist with the systematic review process. These include techniques that identify factors like the presence of randomised control trials within a given study[25], or the specific text sections that correspond to the elements of the PICO[26] framework (population; intervention; comparator; outcomes) [27, 28, 29]. Multiple factors about a paper can be used to build models to filter articles in a systematic review, including the article format, purpose and the scientific rigor of the content [30] or evaluating the quality of the research irrespective of the content[31]. There are multiple datasets available for these tasks[32, 33, 30], and multiple modelling approaches applied to this task with a recent focus on the use of neural networks from Long Short Term Memory (LSTM) [34] to Transformer models[35, 30]. These models can be used either to determine exclusion criteria for filtering abstracts or given to a human reviewer for consideration in the review process.
@@
Review articles from the systematic review literature itself have concluded that machine learning based filters that reduce the volume of articles needed to be reviewed are the most promising line of research in the automation of systematic reviews [36, 37].
Text data has been shown to be useful across a large number of biomedical applications, including the use of unstructured clinical notes in building prognostic models[38]. It has been noted recently that text processing tools are of increasing importance as we attempt to deal with ’infodemics’ of the kind generated by the COVID-19 crisis [39].
Using pure text data models for automating systematic review tasks is also promising. Many of the approaches outlined previously partially use the text data in article titles and abstracts within the models. The long-term goal remains to build models that replicate the evaluation process conducted by human beings, who ultimately look at an article with a research question in mind, and then choose whether to include it in the systematic review.
@@
The Transformer model[40] has proven to be an effective tool for building text-based classification systems across a range of domains. Part of the utility comes from the fact that it can be pre-trained on the language from a domain and then used for specific tasks, allowing the previous learning to be ’transferred’ and applied. It has been shown that transformer models pre-trained on academic abstracts can be used to create state-of-the-art classifiers for systematic review systems, but only when they train the models on data for a specific task [19]. This extends previous work using standard feed-forward neural networks to learn a document embedding that can be employed via transfer learning for a specific systematic review task[16]. The requirement for task-specific training places a burden on teams in terms of the time and human effort required to provide labels for the initial classifier, that is then iteratively improved with more manual labels. The open research question is whether we can use modern natural language processing transformers to develop general-purpose filters that learn the general correspondence between the phrasing of a research question and the relevant articles.
@@
This research outlines a dataset and comparison of methods for a general-purpose systematic review model. A single model is trained that is given input data combining a research question and information about a specific article. It then produces a score indicating the likelihood that the specific article would be included by a human reviewer during a literature pre-screen. Our results demonstrate that using BioBERT[41] version of the BERT Transformer[42] pre-trained on biomedical literature can contribute a promising solution to this problem, but the ideal approach depends on the intended process for using a model to improve the systematic review process. We discuss the observed relationships between research questions and machine learning approaches, and conclude that more work is needed to determine the ideal model for all systematic reviews.
3 Methodology
@@
Figure 1: Schematic representation of the standard machine learning approach to building a systematic review filter tool. 1. Research question is proposed, 2. Database search is conducted, 3. List of candidate articles is compiled, 4. Human annotators review a small selection of articles, 5. Small annotated collection is used to train a machine learning filter, 6. The machine learning is applied to the remaining documents, 7. Producing the final set for complete review by human experts.
We apply machine learning to develop a process for filtering abstracts in a systematic review that is generally applicable and independent of researcher input to provide a question specific label set. This means that it takes information about the specific research question and the content of all candidate abstracts to determine which articles are relevant. This is contrasted with the standard approach to building systematic review tools with machine learning shown in Figure 1.
@@
Figure 2: Schematic representation of the general systematic review filter process. The first five steps define how the general filter is trained, the last two the application to new research questions. 1. A set of training research question are taken, 2. Database searches are conducted for each, 3. Lists of candidate articles are compiled, 4. Human annotators review these training questions to create an annotated set, 5. The general filter is trained. 6. The machine learning is applied to filter documents for a new query, 7. Producing the final set for complete review by human experts.
@@
Our alternative approach is depicted schematically in Figure 2. In the top part of the schematic we see that the model is trained on data derived from multiple research questions that generated human-annotated lists of key articles to review. The bottom frame demonstrates that this general purpose filter can now be applied to future research questions without the need for human annotators to provide an initial seed set of labeled articles for the new research questions. The removal of task-specific annotation for each new research question is the goal of our research; we use machine learning techniques that can understand the match between an outline of a research question and the content of an article’s abstract.
To build this general-purpose filter, we require a dataset of multiple research questions, with all of the articles returned from the database searches, as well as the labels resulting from the initial human screening. The curation of this dataset is described below.
3.1 Dataset Curation
Table 1: Medical Research Questions Datasets
@@ 3.1 Dataset Curation
For the purposes of data curation, we restrict our initial analysis to research questions that are all within the general domain of Therapeutics. This is a constrained domain within the scope of common medical research topics. We identified 10 research questions from the ASERNIP-S systematic review archive for which a team of human researchers had manually labelled thousands of articles as either relevant or not relevant. These initial questions are summarised in Table 1, showing the total number of abstracts retrieved from bibliographic database (e.g, PubMed) searches and the number deemed relevant by the human reviewers.
@@
In addition to the natural language expression of the overall research question, we also generated natural language expressions for each of the four PICO elements independently. The PICO framework has been shown to be well suited to representing therapy questions[43], and we conduct experiments to determine whether text based machine learning tools are able to exploit this alternative framing of the research question.
@@
3.2 Machine Learning Models
Table 2: Text Match Feature Functions
@@ 3.2 Machine Learning Models
We approach this problem using four general machine learning strategies. The first strategy is a baseline approach using a set of standard feature engineering approaches and general-purpose machine learning algorithms ranging from simple to complex. In this approach the research question is contained in a single text field of the data and compares with the article’s title and abstract using text matching algorithms. Each of these text matching approaches generates a numeric value which either describes the context of the research article (title and abstract) or indicates the similarity of the text in the question and the article (title and abstract). The text features and matching approaches are drawn from multiple python packages and compiled into the texturizer python package [44]; the full set of feature generating functions we use are summarised in Table 2, similarly the full set of standard machine learning algorithms are summarised in Table 3. Note, that the TFIDF approach is only used in some of the experiments, and in all cases this is indicated in the descriptions of the results.
@@
Table 3: Machine Learning Models
Model Description Library
Linear Logistic Regression sklearn.linear_model.LogisticRegression
Naive Bayes (GNB) Guassian Naive Bayes sklearn.naive_bayes.GaussianNB
Naive Bayes (CNB) Complement Naive Bayes sklearn.naive_bayes.ComplementNB
SVM Classifier Support Vector Classifier sklearn.svm.LinearSVC
ExtraTrees ExtraTrees Random Forrest sklearn.ensemble.ExtraTreesClassifier
LGBM Light Gradient-Boosting Machine lightgbm.LGBMClassifier
BERT Pre-Trained BERT Transformer transformers.AutoModelForSequenceClassification
In the third approach, we employ a set of pre-trained BERT Transformer models [42] to generate sentence similarity scores as new features. We continue to use the standard text match features from strategies one and two, but simply append the BERT similarity features to the dataset. This approach is done using both the standard natural language representation of the research question and the PICO framework text. We use two general versions of BERT, the BioBERT model[41] and the BlueBERT model[45]. Summary details for these pre-trained BERT models is shown in Table 4, including the number of heads, layers and the vocabulary size of each model. These parameters provide an approximation to the complexity of each model.
@@
In the final fine-tuned transformer approach, we leverage a pre-trained transformer language model to process the data as a question-and-answer model. In this framing, the research question poses a question, the paper abstract provides the context, and the target is the model’s yes or no response. We used the native BERT architecture to accept two input sentences, typically used in pre-training processes with a target value determining if the sentences possess the same meaning. In our usage, this pathway indicates if the abstract (sentence 2) is appropriate for the question (sentence 1).
Table 4: Pre-Trained Transformers
@@
3.3 Experimental Design
We aim to design a system that can identify relevant abstracts based on a new research question proposed for systematic review. The ideal system cannot rely on a bootstrap set of initial labelled examples for the specific research question. To align our testing protocol with this goal we use a variation of the cross-validation methodology in which each fold of the cross-validation consists of all data for one research question. This Leave-One-Question-Out (LOQO) provides a close approximation to real-world usage, we evaluate how well the model can identify the relevant articles based purely on the correspondence between the content of the research question and the content of the article’s title and abstract.
Note that our approach is far more stringent than the standard cross validation methodology. Standard cross validation would involve training and testing on data points from the same research question pool, as such it does not evaluate performance on truly novel research questions.
3.4 Evaluation Metrics
To evaluate our results we use multiple different metrics. The Area Under the Curve ($AUC$)[46] provides a value bounded between zero and one that captures the overall ability of a model to rank all articles. It is calculated by iterating over all possible output thresholds for classification and thereby indicates the quality of the model independently of the classification boundary[47].
In addition, we use two metrics to evaluate the performance of the model at the upper and lower end of the article ranking, each of which cater to specific use cases of the models. We first consider the use case of automatically filtering out articles that are not relevant. Eliminating candidate articles is the dominant method by which machine learning filtering tools can provide efficiency for systematic review processes[36, 37]. To evaluate models for this application we calculate the accuracy in the bottom 50% ($AccBot50\%$). This means we take the 50% of records for a given research question that have the lowest scores. We then calculate the percentage of them that were annotated as irrelevant by the human reviewers. The calculation is shown in Equation 3.1:
@@ 3.4 Evaluation Metrics
We consider a second use case in which a machine learning filter could be used to identify high priority articles to automatically include for full review. To evaluate models for this use case we calculate the accuracy in the top 50 ($AccTop50$). For this metric we take the top 50 highest scoring articles (not percentage) and calculate the accuracy of this set under the assumption that they should all be scored as $relevant$. The calculation is shown in Equation 3.2:
[…]
Where we add an additional function $\mathcal{R}(f(Q_{i}))$ which returns the rank position of the model scored using the function $f$ on article $i$ of question $Q$. This allows us to calculate the score on only the top 50 ranked articles.
4 Results
We show the results for both standard machine learning approaches and transformer models in Tables 5, Table 6 and Table 7. These table display the $AUC$, $AccBot50\%$ and $AccTop50$, respectively. Each modelling approach is displayed as a row with the results for each of the research questions in the corresponding column. We have highlighted in yellow the best (or equal best) score for each research question.
Table 5: AUC by Model and Research Question